# Heart Failure Prediction using Modal Decomposition and Masked Autoencoders for Scarce Echocardiography Databases

### v1

@@ 14 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $104~{}309$
Table 2 shows the experimented cases of the training database $D_{train}$, each one indicating which types of data are used to form that. Different types of test data, resulting from the Modal Decomposition-based Data Transform phase described in Subsection 2.2, have been assessed as well. In this way, the impact of each type of data on the final performance could be evaluated, including the robustness to noise, as the original echocardiography imagery inherently has much noise. Note that the absolute value, the real, and the imaginary parts of the HODMD modes have been taken to increase the number of samples (i.e., images) of the database and so to obtain the total number of training samples in each case. Also note that the final column, i.e., SVD (HODMD Modes), represents the use of the reconstructions of the HODMD modes after applying the SVD algorithm on them. Similarly, the absolute value, and the real and imaginary parts have been taken to increase the database.
@@ 4 Results
The algorithm considered in this paper has been tested with the cardiac database using the different cases for the training database $D_{train}$ presented in Table 2. Additionally, it has been compared with ViT and CNN-based algorithms, including DeiT (Touvron et al. (2021)), ResNet50-version 2 (He et al. (2016)), and Swin-version 2 (Liu et al. (2022)).
@@
$\centering\begin{split}\mu=\mathbb{E}\big{[}\{\hat{P}_{j}\}\big{]}\\ \sigma=\sqrt{\mathbb{E}\big{[}(\{\hat{P}_{j}\}-\mu)^{2}}\big{]}\end{split}\@add@centering$ (4)
@@
$\centering\begin{split}\textit{RMSE}=\sqrt{\mathbb{E}\big{[}(\{\hat{P}_{j}\}-\{T_{j}\})^{2}}\big{]}\\ \textit{Max error}=\max{\big{(}\big{\{}\{\hat{P}_{j}\}-\{T_{j}\}\big{\}}\big{)}}\\ \textit{Min error}=\min{\big{(}\big{\{}\{\hat{P}_{j}\}-\{T_{j}\}\big{\}}\big{)}}\end{split}\@add@centering$ (5)
The maximum and minimum errors have been represented with sign (w/) and without sign (w/), i.e., in absolute value. Therefore, in the case of errors represented with sign, if $\hat{P}_{j}-{T}_{j}<0$, the predicted time would indicate that the heart failure would happen sooner than the real time of heart failure, and, it would be later otherwise. Unless the contrary is explicitly indicated, the heart failure times are represented in months.
The average processing time per sample $\bar{t}$ is the average time taken by the proposed system to process a sample (either image or sequence) and give the heart failure prediction. It is broken down into the SVD algorithm ($\bar{t}_{SVD}$), the HODMD algorithm, comprising the HOSVD dimensionality reduction ($\bar{t}_{HOSVD}$) prior to the HODMD itself ($\bar{t}_{HODMD}$) (Bell-Navas et al. (2023)), and the Deep Neural Network-based Heart Failure Prediction phase ($\bar{t}_{pred}$). Derived from $\bar{t}_{pred}$, the throughput (measured in frames per second, fps) is calculated as the inverse value. Unless the contrary is explicitly specified, the times representing computational cost are measured in milliseconds.
The results have been obtained using a cluster with Intel Xeon Gold 6240R, AMD Ryzen Threadripper PRO 5995WX, three Tesla A100 GPU, and four Nvidia RTX A4500 working in parallel for training, and Intel Xeon Gold 6230 and one Tesla V100 GPU for testing. In this way, not only the high amount of data conforming the created cardiac database is managed, but also the training convergence is sped up and eased.
@@ 4.1 Performance evaluation using different types of data
First, the performance in terms of heart failure prediction accuracy of the training cases on different types of data used for test, i.e., test data obtained with the Modal Decomposition-based Data Transform phase described in Subsection 2.2, is presented. To that end, the configuration of the SVD and HODMD algorithms, the deep neural network, and of the prediction is initially introduced. Later on, the influence of the types of data used for both training and test on the overall performance of the system is studied. Lastly, the heart failure prediction performance in each heart state for the best configuration is shown.
@@
    Configuration: The configuration for the Modal Decomposition-based Data Generation and Transform phases is as follows. For the SVD algorithm, the number of modes provided has been set to $5$, considered adequate for fair reconstructions with the noise filtered according to the contribution of the modes in the frames. Regarding the HODMD algorithm, the number of snapshots used on each sequence $K$ has been set to the total number of frames of the corresponding sequence. In this way, the maximum number of cardiac cycles possible can be captured and the cardiac and respiratory frequencies can be more accurately obtained, and so modes characterizing the physics of the data. The time interval between snapshots $\Delta t=4~{}ms$ and the tolerances $\epsilon_{SVD}=\epsilon_{DMD}=5\mathrm{e}{-4}$ have been selected according to Groun et al. (2022): on the one hand, the value of $\Delta t$ is based on the used ultrasound scanner, configured by a specialist, in the way that the images acquired properly give the cardiac and respiratory frequencies. On the other hand, regarding the selected tolerances for the dimensionality reduction and amplitude truncation steps, respectively, these are larger than the noise level, and lead to a reasonable number of frequencies and so characteristic modes associated with the studied heart states and the heart failures. The index $d$ has been configured for each sequence in the way that fair reconstructions of the snapshots are obtained. In particular, $d=\left\lfloor K/5\right\rfloor$ results adequate for these reconstructions. Precisely, this value of $d$ follows the recommendations for the calibration process described in Le Clainche and Vega (2017). In fact, other values of $d$ lead to reconstructions with severe distortions and artifacts which do not describe the dynamics of the data and do not adequately reproduce the heart cycles. Therefore, discriminative features associated with the heart states and leading to more accurate heart failure predictions are not extracted in these cases.
@@
    Table 3: Comparison results with different types of data for test and training cases using the proposed framework. Refer to Table 2 for the specifications of each training case.
@@
    Performance: Table 3 reveals the impact of the training cases presented in Table 2 using the different types of data for testing on the heart failure time prediction performance of the proposed system. The training cases lead to a general improvement with respect to the case 1 (with only original images), considering that, in general, more training samples are available, as expected. However, the specific types of training data affect the variation of the performance: that is, for example, the case 9, which has more samples than the case 11, leads to a lower performance. The first one uses original images and reconstructions from both the SVD and the HODMD algorithms, but the other incorporates modes instead of the original images. Also note that there is a consistent improvement in the training cases which incorporate HODMD modes, and also in which their SVD reconstructions are included, i.e., cases 7, 10, 11, 12, 13, and 14. Therefore, these two facts reveal the importance of using HODMD modes. This is due to that the HODMD modes describe the dynamics of the data and contain the temporal information useful for the ViT to distinguish the different heart states and so to determine a more accurate heart failure time. In addition, the data generated with the SVD and the HODMD algorithms have less noise than the original data (inherently with much noise), once more, leading to more discriminative features. This shows the importance of using the SVD and the HODMD algorithms for both data augmentation and extraction of more discriminative features associated to the different heart states, allowing more accurate heart failure prediction times. Therefore, these algorithms improve the discriminating capability of the proposed ViT in the usual scenario of having a scarce number of samples. In addition, this also demonstrates the usefulness of the modal decomposition techniques to deal with the difficulty to elaborate varied databases with high quality enough in the medical field in both costs and very hard specialized work. On the other hand, the results show the effectiveness of the SVD and of the HODMD algorithms to deal with the high amount of noise inherent to the original echocardiography images, leading to performance improvements, and making the proposed system more robust to noise. If another type of noise were added to the echocardiography images, the SVD and the HODMD algorithms would also effectively clean the data and would generate more discriminative features which lead to more accurate heart failure time predictions.
@@
    Table 4 presents the heart failure prediction results for each heart state using the best configuration inferred from Table 3: training case 14 and original images as test data. This kind of test data has been selected because of having more original images than modes, leading to more reliable prediction results in the fusion of the estimated heart failure times. The performance is balanced among the considered heart states, like the number of test samples in terms of both images and sequences. A general error margin of $4.54$ months is reached, which is high enough for medical applications considering the low number of original training samples (images) for each class, much lower than the samples required by typical deep learning algorithms (hundreds of thousands or even millions of samples). This means that, if a much higher number of samples were available, the heart failure times would be more accurately predicted and the features associated to the different heart states better learned. However, this is unfeasible, considering the high difficulty in terms of costs and specialized hard work in the medical field. On the other hand, according to the means and standard deviations, the proposed system can estimate heart failure times similar to the real values and follow the general statistics. Therefore, the results obtained demonstrate the capability of the proposed ViT to effectively model the non-linear relationship which could be present between the input and the output: in this case, between echocardiography images and heart failure times.
@@ 4.2 Comparison with Alternative Algorithms
In this second part of the experiments, the performance of the proposed system has been compared against other alternative deep neural networks. This is, using the same framework, the proposed ViT has been replaced by other deep neural networks. In particular, the compared deep neural networks are DeiT (Touvron et al. (2021)), ResNet50-version 2 (He et al. (2016)), and Swin-version 2 (Liu et al. (2022)). First, the configuration of the proposed system is introduced. Next, the configuration and the necessary adaptations of the alternative algorithms are described. Finally, an overall performance comparison between the proposed ViT and the other deep neural networks is presented, using the most representative training cases (1, 7, 10, 11, 12, 13, 14) and the main types of test data (Original, HODMD-based reconstructions, and HODMD modes).
Table 4: Heart failure time prediction performance using $\mu$, $\sigma$, error margin, maximum and minimum errors (with and without sign) obtained with the best configuration of the proposed framework.
Heart State Age of heart failure Min error # Test samples
Real Predicted Error margin Max error (with sign) (w/) (w/o) Images Sequences
@@
    Performance: Table 5 presents a comparison of the heart failure time prediction performance, using the cardiac database and the proposed system with either the ViT trained with the proposed self-supervised learning approach, DeiT-S, Swin-v2-T, or ResNet50-v2, with pretraining or with no pretraining. Only the training cases 1, 7, 10, 11, 12, 13, and 14 have been considered, as being the most representative ones regarding the contribution of the SVD and HODMD algorithms with respect to the use of only original images. Regarding the types of test data, original images, HODMD-based reconstructions and HODMD modes have been compared, as mainly representing the influence of the SVD and HODMD algorithms as feature extractors.
    As can be observed in the table, the two best results in terms of heart failure prediction performance have been obtained using the proposed ViT and the Swin-v2-T, according to the error margins represented. Moreover, these algorithms even outperform ResNet50-v2, showing the efficacy of using more complex models, consisting in Transformers which incorporate mechanisms to better train them and adapt them more to scarce datasets (e.g., the LIFE module to increase the receptive field in the inputs of the attention mechanisms in the case of the Swin Transformer, or the self-supervised learning approach proposed for the ViT). As deep neural networks, both show that they can effectively model the non-linear relationships which could be present between echocardiography imagery and heart failure times. On the other hand, there is a general improvement in the training cases with respect to the case 1 and when using the HODMD modes for test data instead of original images. Therefore, the alternative algorithms can also benefit from the use of the SVD and HODMD algorithms for both data augmentation and feature extraction, also incorporating temporal information thanks to the HODMD modes. Once more, this improvement also demonstrates the effectiveness to address the high noise inherent to the original echocardiography images and, therefore, the robustness of the proposed system with the alternative algorithms to noise. Again, other kinds of noise additionally added to the echocardiography imagery would be effectively reduced by the SVD and HODMD algorithms, leading to performance improvements with the alternative algorithms. On the other hand, note that, in the training cases without SVD nor HODMD modes, like the case 1, a better performance is generally achieved when using original test images instead of modes. This can be attributed to the fact that the models learn to predict heart failures in original images instead of in images represented with the most essential features, obtaining more confident predictions in the original images.
@@
    Regarding the results obtained, the training cases with HODMD-based data has led to the best performances among the compared algorithms. In the specific case of the proposed ViT, the training case 14 gives the best results, showing the contribution of all types of data. Then, in general, the use of original images or SVD-based data for training can also increase, although with lower difference, the final performance, as can be seen in the case of the Swin-v2-T.
    Table 6 summarizes the computational cost of the different phases of the prediction process. Note that the SVD and the HODMD algorithms are applied on sequences rather than on single images. However, the time measures represented in the table are averages per image. As can be seen, the HOSVD dimensionality reduction process implies the highest computational cost by far (Bell-Navas et al. (2023)). On the other hand, the proposed ViT trained with self-supervised learning is practically the fastest deep learning algorithm, which is desirable for several real-time medical applications. This has little difference with respect to ResNet50-v2, but the proposed ViT achieves a better trade-off between computational cost and prediction performance. In addition, the proposed ViT has the lowest number of learnable parameters among the compared algorithms. Therefore, these results encourage the deployment of the proposed ViT using less resources. In every case, all time values are inside of real-time requirements, which is appealing for several applications in the medical field, for example, in portable devices, allowing to cover more remote areas.
@@
    Table 5: Comparison of the recognition performance among different algorithms with the main types of test data and the main training cases. Refer to Table 2 for the specifications of each training case.

### v2

@@ 14 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $104~{}309$
Table 2 shows the experimented cases of the training database $\boldsymbol{D}_{train}$, each one indicating which types of data are used to form that. Different types of test data, resulting from the Modal Decomposition-based Data Transform phase described in Subsection 2.2, have also been assessed. In this way, the impact of each type of data on the final performance could be evaluated, including the robustness to noise, as the original echocardiography imaging inherently has much noise. Note that the absolute value, the real, and the imaginary parts of the HODMD modes have been taken to increase the number of samples (i.e., images) of the database and so to obtain the total number of training samples in each case. Also note that the final column, i.e., SVD (HODMD Modes), represents the use of the reconstructions of the HODMD modes after applying the SVD algorithm on them, that is, $\boldsymbol{D}_{SVD2\_R}$, according to the notation presented in Subsection 2.1. Similarly, the absolute value, and the real and imaginary parts have been taken to increase the database.
@@ 4 Results
The algorithm considered in this paper has been tested with the cardiac database using the different cases for the training database $\boldsymbol{D}_{train}$ presented in Table 2. Additionally, it has been compared with ViT and CNN-based algorithms, including DeiT (Touvron et al. (2021)), ResNet50-version 2 (He et al. (2016)), and Swin-version 2 (Liu et al. (2022)).
@@
$\centering\begin{split}\mu=\mathbb{E}\big{[}\{\hat{P}_{j}\}\big{]}.\\ \sigma=\sqrt{\mathbb{E}\big{[}(\{\hat{P}_{j}\}-\mu)^{2}}\big{]}.\end{split}\@add@centering$ (4)
@@
$\centering\begin{split}\textit{RMSE}=\sqrt{\mathbb{E}\big{[}(\{\hat{P}_{j}\}-\{T_{j}\})^{2}}\big{]}.\\ \textit{Max error}=\max{\big{(}\big{\{}\{\hat{P}_{j}\}-\{T_{j}\}\big{\}}\big{)}}.\\ \textit{Min error}=\min{\big{(}\big{\{}\{\hat{P}_{j}\}-\{T_{j}\}\big{\}}\big{)}}.\end{split}\@add@centering$ (5)
The maximum and minimum errors have been represented with and without sign (that is, in absolute value). For brevity, error with sign is denoted as (w/), and error without sign as (w/o). Therefore, in the case of errors represented with sign, if $\hat{P}_{j}-{T}_{j}<0$, the predicted time would indicate that the heart failure would happen sooner than the real time of heart failure, and it would be later otherwise. Unless the contrary is explicitly indicated, the heart failure times are represented in months.
The average processing time per sample $\bar{t}$ is the average time taken by the proposed system to process a sample (either image or sequence) and give the heart failure prediction. It is broken down into the SVD algorithm ($\bar{t}_{SVD}$), the HODMD algorithm, comprising the HOSVD dimensionality reduction ($\bar{t}_{HOSVD}$) prior to the HODMD itself ($\bar{t}_{HODMD}$) (Bell-Navas et al. (2023)), and the Deep Neural Network-based Heart Failure Prediction phase ($\bar{t}_{pred}$). The throughput is measured in frames per second, fps, and represents the number of snapshots which can be processed by the deep neural network in a second. Therefore, this metric assesses the capability of the deep neural network for real-time applications. The throughput is calculated as $1/\bar{t}_{pred}$. Unless the contrary is explicitly specified, the times representing computational cost are measured in milliseconds.
The results have been obtained using a cluster with Intel Xeon Gold 6240R, AMD Ryzen Threadripper PRO 5995WX, three Tesla A100 GPU, and four Nvidia RTX A4500 working in parallel for training, and Intel Xeon Gold 6230 and one Tesla V100 GPU for testing. In this way, not only the high amount of data conforming the created cardiac database is managed, but also the training convergence is speeded up and eased.
@@ 4.1 Performance evaluation using different types of data
First, the performance in terms of heart failure prediction accuracy on the training cases and on different types of data used for test, i.e., test data obtained with the Modal Decomposition-based Data Transform phase described in Subsection 2.2, is presented. To that end, the configuration of the SVD and HODMD algorithms, the deep neural network, and of the prediction is initially introduced. Later on, the influence of the types of data used for both training and test on the overall performance of the system is studied. Lastly, the heart failure prediction performance in each heart state for the best configuration is shown.
@@
    Configuration: The configuration for the Modal Decomposition-based Data Generation and Transform phases is as follows. For the SVD algorithm, the number of modes provided has been set to $5$, considered adequate for fair reconstructions with the noise filtered according to the contribution of the modes in the frames. Regarding the HODMD algorithm, the minimum number of snapshots a sequence must have to be able to apply the HODMD algorithm has been set to $100$, following the practices of Groun et al. (2022). In this way, at least three heart cycles are covered, considered enough to obtain fair reconstructions and to correctly characterize the temporal information. Moreover, the number of snapshots taken from each sequence $K$ has been set to the total number of frames of the corresponding sequence. In this way, the maximum number of cardiac cycles possible can be captured and the cardiac and respiratory frequencies can be more accurately obtained, and so modes characterizing the physics of the data. The time interval between snapshots $\Delta t=4~{}ms$ and the tolerances $\epsilon_{SVD}=\epsilon_{DMD}=5\mathrm{e}{-4}$ have been selected according to Groun et al. (2022): on the one hand, the value of $\Delta t$ is based on the used ultrasound scanner, configured by a specialist, in the way that the images acquired properly give the cardiac and respiratory frequencies. On the other hand, regarding the selected tolerances for the dimensionality reduction and amplitude truncation steps, respectively, these are larger than the noise level, and lead to a reasonable number of frequencies and so characteristic modes associated with the studied heart states and the heart failures. The index $d$ has been configured for each sequence in the way that fair reconstructions of the snapshots are obtained. In particular, $d=\left\lfloor K/5\right\rfloor$ results adequate for these reconstructions. Precisely, this value of $d$ follows the recommendations for the calibration process described in Le Clainche and Vega (2017). In fact, other values of $d$ lead to reconstructions with severe distortions and artifacts which do not describe the dynamics of the data and do not adequately reproduce the heart cycles. Therefore, discriminative features associated with the heart states and leading to more accurate heart failure predictions are not extracted in these cases.
@@
    Table 3: Error margin results comparing different types of data for test and training cases using the proposed framework (best results in bold, in Case 14). Refer to Table 2 for the specifications of each training case.
@@
    Performance: Table 3 reveals the impact of using the different types of data for testing and of the training cases presented in Table 2 on the heart failure time prediction performance of the proposed system. In this table, the error margin has been represented. The training cases lead to a general improvement, because of the decreasing error margin values, with respect to the case 1 (with only original images), considering that, in general, more training samples are available, as expected. Precisely, the case 14 which considers all types of data (original images and both SVD and HODMD modes and reconstructions) leads to the best error margin: $4.54$ months. On the other hand, in this case 14, the performance barely changes, regardless of the type of testing data used, which demonstrates the robustness of the proposed system to noise and to different types of data. On the other hand, the training cases affect the variation of the performance: for example, the case 9, which has more samples than the case 11, leads to a lower performance. The case 9 uses original images and reconstructions from both the SVD and the HODMD algorithms, but the other one, the case 11, incorporates modes instead of the original images. Also note that there is a consistent improvement in the training cases which incorporate HODMD modes, and also in those training cases which include the SVD reconstructions of the HODMD modes, which are the cases 7, 10, 11, 12, 13, and 14. Therefore, these two facts reveal the importance of using HODMD modes. This is due to that the HODMD modes describe the dynamics of the data and contain the temporal information useful for the ViT to distinguish the different heart states and so to determine a more accurate heart failure time. In addition, the data generated with the SVD and the HODMD algorithms have less noise than the original data (inherently with much noise), once more, leading to more discriminative features. This shows the importance of using the SVD and the HODMD algorithms for both data augmentation and extraction of more discriminative features associated to the different heart states, allowing more accurate heart failure prediction times. Therefore, these algorithms improve the discriminating capability of the proposed ViT in the usual scenario of having a scarce number of samples. In addition, this also demonstrates the usefulness of the modal decomposition techniques to deal with the difficulty to elaborate varied databases with high quality enough in the medical field in both costs and very hard specialized work. On the other hand, the results show the effectiveness of the SVD and of the HODMD algorithms to deal with the high amount of noise inherent to the original echocardiography images, leading to performance improvements, and making the proposed system more robust to noise. If another type of noise were added to the echocardiography images, the SVD and the HODMD algorithms would also effectively clean the data and would generate more discriminative features which lead to more accurate heart failure time predictions.
@@
    Table 4 presents the heart failure prediction results for each heart state using the best configuration inferred from Table 3: training case 14 and original images as test data. This kind of test data has been selected because of having more original images than modes, leading to more reliable prediction results in the fusion of the estimated heart failure times. The performance is balanced among the considered heart states, like the number of test samples in terms of both images and sequences. A general error margin of $4.54$ months is reached, which is accurate enough for medical applications considering the low number of original training samples (images) for each class, much lower than the samples required by typical deep learning algorithms (hundreds of thousands or even millions of samples). This means that, if a much higher number of samples were available, the heart failure times would be more accurately predicted and the features associated to the different heart states better learned. However, this is unfeasible, considering the high difficulty in terms of costs and specialized hard work in the medical field. On the other hand, the means ($\mu$) and standard deviations ($\sigma$) of the real and predicted ages of heart failures are similar, which means that the proposed system can follow the statistics of the real ages of heart failures under the different heart states, and therefore estimate heart failure times close to the real ones. Therefore, the results obtained demonstrate the capability of the proposed ViT to effectively model the non-linear relationship which could be present between the input and the output: in this case, between echocardiography images and heart failure times.
@@ 4.2 Comparison with Alternative Algorithms
In this second part of the experiments, the performance of the proposed system has been compared against other alternative deep neural networks. This is, using the same framework, the proposed ViT has been replaced by other deep neural networks. In particular, the compared deep neural networks are DeiT (Touvron et al. (2021)), ResNet50-version 2 (He et al. (2016)), and Swin-version 2 (Liu et al. (2022)). These deep neural networks have been selected for the comparisons to cover a range of competitive CNNs and ViTs which introduce alternative mechanisms for better training regarding the architecture and adoption of auxiliary databases. First, the configuration of the proposed system is introduced. Next, the configuration and the necessary adaptations of the alternative algorithms are described. Finally, an overall performance comparison between the proposed ViT and the other deep neural networks is presented, using the most representative training cases (1, 7, 10, 11, 12, 13, 14) and the main types of test data (Original, HODMD-based reconstructions, and HODMD modes).
Table 4: Heart failure time prediction performance for the studied heart states using $\mu$, $\sigma$, error margin, maximum and minimum errors (with: w/, and without sign: w/o) obtained with the best configuration of the proposed framework (training case 14 and original images as test data. Refer to Table 2 for the specifications of each training case).
Heart State Age of heart failure ($\mu\pm\sigma$) Min error # Test samples
Real Predicted Error margin Max error (w/) (w/) (w/o) Images Sequences
@@
    Performance: Table 5 presents a comparison of the heart failure time prediction performance in terms of the error margin, using the cardiac database and the proposed system with either the ViT trained with the proposed self-supervised learning approach, DeiT-S, Swin-v2-T, or ResNet50-v2, with pretraining or with no pretraining. Only the training cases 1, 7, 10, 11, 12, 13, and 14 from Table 2 have been considered, as being the most representative ones regarding the contribution of the SVD and HODMD algorithms with respect to the use of only original images. Regarding the types of test data, original images, HODMD-based reconstructions and HODMD modes have been compared, as mainly representing the influence of the SVD and HODMD algorithms as feature extractors.
    As can be observed in the table, the two best results in terms of heart failure prediction performance have been obtained using the proposed ViT and the Swin-v2-T, according to the error margins represented. Moreover, these algorithms even outperform ResNet50-v2, showing the efficacy of using more complex models, consisting in Transformers which incorporate mechanisms to better train them and adapt them more to scarce datasets (e.g., the LIFE module to increase the receptive field in the inputs of the attention mechanisms in the case of the Swin Transformer, or the self-supervised learning approach proposed for the ViT). As deep neural networks, both show that they can effectively model the non-linear relationships which could be present between echocardiography imaging and heart failure times. On the other hand, there is a general improvement in the training cases with respect to the case 1 and when using the HODMD modes for test data instead of original images. Therefore, the alternative algorithms can also benefit from the use of the SVD and HODMD algorithms for both data augmentation and feature extraction, also incorporating temporal information thanks to the HODMD modes. Once more, this improvement also demonstrates the effectiveness to address the high noise inherent to the original echocardiography images and, therefore, the robustness of the proposed system with the alternative algorithms to noise. Again, other kinds of noise additionally added to the echocardiography imaging would be effectively reduced by the SVD and HODMD algorithms, leading to performance improvements with the alternative algorithms. On the other hand, note that, in the training cases without SVD nor HODMD modes, like the case 1, a better performance is generally achieved when using original test images instead of modes. This can be attributed to the fact that the models learn to predict heart failures in original images, obtaining more confident predictions in the original images than in other representations.
@@
    Regarding the results obtained, the training cases with HODMD-based data have led to the best performances among the compared algorithms. In the specific case of the proposed ViT, the training case 14 gives the best results (error margin of $4.54$ months), showing the contribution of all types of data. In general, the use of original images or SVD-based data for training can also increase the final performance, although with a lower difference, as can be seen in the case of Swin-v2-T.
    Table 6 summarizes the computational cost of the different phases of the prediction process. Note that the time measures represented in the table are averages per image, although the SVD and the HODMD algorithms are applied on sequences rather than on single images. As can be seen, the HOSVD dimensionality reduction process implies the highest computational cost by far (Bell-Navas et al. (2023)). On the other hand, the proposed ViT trained with self-supervised learning is practically the fastest deep learning algorithm, with a much higher throughput than the one of Swin-v2-T, which is desirable for several real-time medical applications. The proposed ViT has a similar $\bar{t}_{pred}$ than the one obtained with ResNet50-v2, but the proposed ViT achieves a better trade-off between computational cost and prediction performance. In addition, the proposed ViT has the lowest number of learnable parameters among the compared algorithms. Therefore, these results encourage the deployment of the proposed ViT using less resources. In every case, all time values are inside of real-time requirements, which is appealing for several applications in the medical field, for example, in portable devices, allowing to cover more remote areas.
@@
    Table 5: Comparison of the heart failure time prediction performance in terms of the error margin among different deep neural networks using the main types of test data and the main training cases (best results in bold, in Case 14). Refer to Table 2 for the specifications of each training case.